A coin lying on a meterstick starts to slide down when one end of the stick is raised 36.4cm above the horizontal. Find The limiting angle for friction and the coefficient of starting friction

1
Expert's answer
2022-04-26T14:56:14-0400

a)

"\theta=\arctan\frac lm=20°,"

b)

"\mu=\frac lm=0.364."